DIATOM - DIATOM POTENTIAL ENERGY SURFACE FOR THE INTERACTION BETWEEN CN(A^{2}Π) AND H_{2}

\underline{U. SCHNUPF}, Department of Chemistry, Troy State University, Troy, Alabama 36082; K. MOROKUMA AND M. C. HEAVEN, Department of Chemistry and Cherry L. Emerson Center for Scientific Computation, Emory University, Atlanta, Georgia 30322.

There has been considerable recent interest in van der Waals complexes formed between open and closed shell diatoms. In the case of CN-H_{2} spectra have been recorded for the B^{2}Σ-X^{2}Σ and A^{2}Π-X^{2}Σ transitions by Chen et al.^{a} Previous calculations show that CN+H_{2} in its electronic ground state has a collinear equilibrium structure^{b} and there is barrier to the reaction H_{2}+CN⟶H+HCN proceeding via a linear transition state. Until now no similar information was available for the CN-H_{2} in its electronic excited state. The purpose of our investigation was to generate an accurate Ab Initio potential energy surface for the CN(A^{2}Π)+H_{2} in the Van der Waals interaction region. A four-dimensional Ab Initio potential energy surface will be reported for the long-range interaction of H_{2} with the open-shell CN diatomic in the its first electronic state A^{2}Π. The potential energy surface is based on a second order multireference perturbation calculation (CASPT2) utilizing a large correlation consistent basis set. The diatomic bond lengths were held constant in these calculations. The properties of the potential surface are compared with the spectrum reported earlier by Chen et al.^{a} In addition, the potential energy surface is compared to recently reported Ab Initio calculations for H_{2}+O_{2}.

aY. Chen and M. C. Heaven J. Chem. Phys. \underline{109}, 5171(1998).

bA. L. Kaledin, M. C. Heaven, and J. M. Bowman J. Chem. Phys. \underline{110}, 10380(1999).